Exercise 2-13 DepartmentalP Wh ite Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates: Direct labor-hours Machine-hours Total fixed manufacturing overhead cost Variable manufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour Cutting Finishing 30,8e0 5,000 264,000 $ 366,08e 6,000 48,000 $ 2.e0 $ 4.00 Required: 1. Compute the predetermined overhead rate for each department 2. The job cost sheet for Job 203, which was started and completed during the year, showed the following Department Cutting Finishing Direct 1labor-hours Machine-hours Direct materials Direct labor cost 20 80 $ 500$310 108 $ 360 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203 3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates? Complete the question by entering your answers in the tabs given below
Answer #1
Answer 1)
Cutting Finishing
Total fixed manufacturing overhead cost = 264000 366000
Total machine hr. 48000
Total labor hr 5000
Total fixed overhead per machine hr =264000/48000 5.50
Total fixed overhead per Labor hr =366000/5000 73.20
Variable manufacturing overhead rate per MH 2
Variable manufacturing overhead rate per labor hr 4
Total overhead rate per machine hr 7.50
Total overhead rate per labor hr 77.20
answer 2) Computation of total manufacturing cost for Job 203 Job 203
Cutting Finishing Total
A Direct material 500 310 810
B Direct labor 108 360 468
Manufacturing overhead
C Cutting =7.5*80 600 600
D Finishing =77.2*4 308.8 308.8
E=A+B+C+D Total manufacturing cost 1208 978.8 2186.8
Therefore total cost = 2186.80
answer 3)
Computation of plant wide overhead rate = =(264000+366000)/(6000+30000) 17.50
Computation of total cost of job
A Direct material 810
B Direct labor 468
C Overhead = 17.5*(6+20) 455
Total cost 1733
Therefore we are seeing cost is lower and is significantly different